OpenStudy (anonymous):

find the area buonded by the curve ay= xsquare and the lines y=a and y=4a

OpenStudy (anonymous):

 \[\int\limits_{0}^{a^4}\int\limits_{a}^{4a}dydx + \int\limits_{a^4}^{4a^4}\int\limits_{\sqrt{x}/a}^{4a}dydx\]